OpenStudy (anonymous):

If f(x,y) = x(raised to 3) + y (raised to 3) --------------------------- x-y ; where x=/= y, and f(x,y) = 0; when x=y. Then find partial derivatives with respect to x (i.e. df/dx )and with respect to y.. (i.e. df/dy) I'm stumped at this question. Can someone help me with this?

OpenStudy (anonymous):

first you must find out whether it is differentiable at x=y

OpenStudy (anonymous):

no you take partial derivatives

OpenStudy (anonymous):

basically it means when respect to x, all y's are constants

OpenStudy (anonymous):

and vice versa

OpenStudy (anonymous):

i have a feeling limitless is doing this he's been typing for quite a while

OpenStudy (anonymous):

@Outkast3r09, are you certain this isn't just easy partial differentiation? I think one has to apply the product rule and that's the hardest part. Besides that, you separate constants and apply a substitution.

OpenStudy (anonymous):

So u would first take a partial derivative with respect to x keeping y constant and then with repect to y keeping x constant. and yes I hope Limitless is solving. :) - Thats wat i feel too.. :3

OpenStudy (anonymous):

Nope, I was solving. I saw that, based on my understanding of partials, it was fairly easy if you knew the right path.

OpenStudy (anonymous):

well , f(x,y)=0 might not be differentiable i'm not sure off the top of my head as this was right out of my head at the end of calc 3

OpenStudy (anonymous):

If \(f(x,y)=0,\) and it is not differentiable at \(x=y\), wouldn't you simply state that in your solution?

OpenStudy (anonymous):

yes however you have to check to make sure

OpenStudy (anonymous):

this is one of thos problems where you have topick specific paths to find if a limit exists

OpenStudy (anonymous):

@Outkast3r09, I am not as familiar with the conditions of differentiation. Couldn't you simply say, "At \(x=y\), the function is undefined. Ergo, it is not differentiable there."

OpenStudy (anonymous):

^Specific paths?????

OpenStudy (anonymous):

i've found a problem much like this let me show you

OpenStudy (anonymous):

kkk plz. :)

OpenStudy (anonymous):

it must be continous to differentiate

OpenStudy (anonymous):

once you've shown that it is not differentiable all you have to do is show that when x doesn't equal y your answer will be the partials

OpenStudy (anonymous):

Wait how do u get -3x^(2)/2x^(2) in the question u gave??

OpenStudy (anonymous):

So, take left and right hand limits. Presto.

OpenStudy (anonymous):

they approached it from y=x so instead of as x,y)-> (0,0) it was as (x,x)-> 0 @dpaInc some assistance please

OpenStudy (anonymous):

the denominator has an x, you have to use product rule and you might have to use an y=f(x) sub

OpenStudy (anonymous):

\(\partial\) (\partial) is used for partial derivatives, not \(d.\)

OpenStudy (anonymous):

^Yyeah i know. Dunno how to insert with keyboard. I assume u added through the equation button??

OpenStudy (anonymous):

anyways, f(x,y)=0 is differential i believe as going from any line will get you 0

OpenStudy (anonymous):

and since it simply says to take the derivatives, it's obvious that the first part of piecewise is also continuous however you' have to check it to make sure

OpenStudy (anonymous):

if you didn't ask

OpenStudy (anonymous):

Can i get something straight???? (Yes I'm terrible @derivatives/integration etc). If u solve delta f/delta x = x^3+y^3/x-y wouldn't u do it using the U/v methood????

OpenStudy (anonymous):

you could or you can do it by product rule \[\frac{\delta f}{dx}=[(x^3+y^3)(x-y)^{-1}]'_x\]

OpenStudy (anonymous):

\[\begin{align} \frac{\partial f}{\partial x}&=\frac{\partial}{\partial x}\frac{x^3}{x-y}+\frac{\partial}{\partial x}\frac{y^3}{x-y}\\ &=\underbrace{\frac{\partial}{\partial x}\frac{x^3}{x-y}}_{\text{product rule}}+y^3\underbrace{\frac{\partial}{\partial x}\frac{1}{x-y}}_{u \text{ sub}} \end{align}\] Can you work from here?

OpenStudy (anonymous):

hence, = [(derivative of u)v-(derivative of v)u]/v^2??? In this case => =[(3x^2)(x-y)] - [(1)(x^3+y^3)[/ (x-y)^2

OpenStudy (eyust707):

\[f(x,y) = {x^3 + y^3 \over x - y} = (x^3 + y^3 )(x-y)^{-1}\] \[{\partial f \over \partial x} = (x^3 + y^3) (-x+y)^{-2} + (x-y)^{-1} (3x)\] \[{\partial f \over \partial y} = (x^3 + y^3) (x-y)^{-2} + (x-y)^{-1} (3y)\]

OpenStudy (anonymous):

I can't remember, for partials, do you need to make a sub to get rid of y's or was that something else

OpenStudy (eyust707):

with partials all other variables are constant

OpenStudy (anonymous):

I'm really sorry, I dunno how to use the euqation thing. :/ So it may be very hard for u to understand wat I'm typing. :(

OpenStudy (anonymous):

i mean't at the product i don't think it's needed , anyways eyust answer is corect you just need to make a common denominator

OpenStudy (anonymous):

I remember something in calculus 3 you had to change everything... i think that was when you had a parameter and you had to switch it all out of x and y's to s and t's

OpenStudy (anonymous):

I'm pretty sure its not the switching both x and y. to s and t. nor to r and :theta:

OpenStudy (anonymous):

o no not this question i'm just saying that later on you'll do something like that and it's insanely annoying. anyways did you make a common denominator?

OpenStudy (anonymous):

@eyust707 thanx for the wolfram links. Seems really helpful. Also Outkast I really appreciate ur help. :) I think I'll be ab;le to easily figure out from the link information. :D thank u again guys. (ALso @Limitless)

OpenStudy (anonymous):

You're welcome. :D

OpenStudy (anonymous):

:) I guess I close this thread now. Thanx to u folks. :)